# Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model

### v2

@@ 5 Related Work
With the rapid development and widespread deployment of LRMs, many works have paid attention to security of LRMs. Several studies  Jiang et al. (2025); Parmar and Govindarajulu (2025) have conducted comprehensive security evaluations on LRMs, e.g., DeepSeek-R1 Guo et al. (2025), revealing existing vulnerabilities of these models. In addition, Zhou et al. (2025) introduce a CoT jailbreak attack method, specifically targeting reasoning models. On the other hand, to mitigate the issues of security, recent studies Jiang et al. (2025); Zhang et al. (2025b) have constructed the safety data with the thought process for supervised fine-tuning of LRMs. However, the above-mentioned works are all confined to the text-only models. Although SafeMLRM Fang et al. (2025) explores the security of MLRMs, the limitation of this work is that it only focuses on one security scenario, i.e., jailbreak robustness. The introduction of another modality in LRMs brings about significant shifts in model behavior, making it crucial to systematically assess the security challenges of MLRMs and develop effective defense methods.
@@
Current studies enhancing the safety capabilities of MLLMs can be categorized into two types: the training-based method and the training-free method. The training-based method typically includes Supervised Fine-Tuning (SFT) Zong et al. (2024); Ding et al. (2025b); Li et al. (2024) and Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2024). In addition, Lee et al. (2024) introduce the weight merging approach to mitigate safety degradation. Another branch of studies incorporates additional safety components in a training-free manner. For instance, MLLM-Protector Pi et al. (2024) offers a plug-and-play detector for harmful responses, and ETA Ding et al. (2025a) proposes an inference-time alignment framework to ensure safety compliance during deployment. In contrast to these methods, our study is the first to investigate the safety behaviors of reasoning models in both jailbreak and awareness of safety scenarios.
@@ A.1 Benchmark and Evaluation Details
For the MM-SafetyBench, FigStep and SIUO, we employ GPT-4o-mini Hurst et al. (2024) and adopt identical configurations in teir work to ensure consistency across assessments. For the JailBreaKV, we follow the same evaluation settings of their work and utilize the Llama Guard 3 Vision Chi et al. (2024) as the evaluator. However, similar to the issues observed in MIS Ding et al. (2025b), we identify shortcomings in the safe category evaluation setup of MSSBench. As a result, we employed human evaluation as an alternative. Additional details are provided in Appendix A.2. The benchmarks are listed as follows:
@@
JailBreaKV Luo et al. (2024) is designed to evaluate the transferability of jailbreak techniques from LLMs to MLLMs, encompassing both text-based and image-based jailbreak attacks. Following the evaluation protocol of VLGuard, we conduct our experiments on the miniset provided in their paper, which contains 280 samples, following the setting of Hu et al. (2024).
@@ A.2 MSSBench Evaluation Details
In the evaluation setup of MSSBench, LLM are required to classify responses into four safety levels. For certain examples in the safe category, responses that provide helpful advice but also emphasizing potential risks are evaluated as incorrect. As shown in Figure 6, the model provides the suggestion of using a heavy-duty drill while emphasizes safety as a priority. To address this limitation, we emplyed human evaluation as an alternative.
@@ A.3 Model Details
We evaluate a total of 11 MLRMs, Table 6 summarizes the Multimodal Large Reasoning Models evaluated for safety and their corresponding base models.
@@ B.1 Detailed Pipeline
To accommodate the input requirements of DeepSeek-R1 Guo et al. (2025), we convert images into caption form as input in advance.In particular, we employ Qwen2.5-VL-72B Bai et al. (2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of security issues, with the prompt “Please provide a detailed description of this image.”.
@@
The primary objective at this step is to filter and transform the original thought process. In particular, the caption field is processed to prevent potential misinterpretation by the MLRM. The original thought process from the previous step also contain redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted. Specific prompts used for this step are provided in the Figure 7.
Model Chat Embodied AVG.
Safe Unsafe AVG. Safe Unsafe AVG.
Gemini2.0-Flash-Thinking 85.67 45.00 65.33 93.06 5.56 49.31 57.32
Claude3.7-Sonnet-Thinking 75.84 45.97 60.91 93.42 15.79 54.61 57.76
QVQ-Preview 80.67 24.67 52.67 53.95 43.42 48.68 50.68
Skywork-R1V 86.67 15.00 50.83 98.68 1.32 50.00 50.42
Llama 3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
LlamaV-o1 59.33 47.00 53.17 100.00 0.00 50.00 51.59
Llama 3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
LLaVA-CoT 94.33 12.33 53.33 100.00 0.00 50.00 51.67
Llama 3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
Mulberry-Llama 81.33 35.00 58.17 97.37 2.63 50.00 54.09
Qwen2.5-VL-3B(base) 83.67 31.00 57.33 59.21 35.53 47.37 52.35
LMM-R1 83.00 33.00 58.00 69.70 26.32 48.03 53.02
Qwen2.5-VL-7B(base) 93.33 10.67 52.00 93.42 2.63 48.03 50.02
R1-Onevision 86.60 15.12 50.86 86.57 7.46 47.01 48.94
InternVL-2.5-8B (base) 92.00 11.33 51.67 98.68 1.32 50.00 50.84
$M$-Eureka 90.67 11.67 51.17 97.37 2.63 50.00 50.59
Kimi-VL (base) 91.67 8.67 50.17 100.00 1.43 50.71 50.44
Kimi-VL-Thinking 83.33 23.67 53.5 98.67 0.00 49.33 51.42
@@ B.3 Prompt for Data Construction
This section presents the prompt and safety regulation used for safety thought process collection. We referenced the definition of User Requests Categorization from Wang et al. (2024) and listed all prompts used in the data construction in Figure 7. During the collection of safety thought process data, we supplied the model with category-specific safety regulations derived from the raw data classification. This strategy not only reduce the length of prompt but also enable tiered defensive measures for distinct safety-related issues.
@@ C.2 Traning Details
For the training procedure, we utilize the LoRA Hu et al. (2022) to fine-tune the R1-Onevision and LLaVA-CoT based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework. For SPA-VL dataset, we use the DPO script for training. All training experiments are conducted using one NVIDIA A100-80G GPUs. The hyper-parameters for supervised fine-tuning are listed in Table 9.
@@ D.2 Performance of Thought Process on MM-SafetyBench
Safe Thought 12.65% 1.90%
Unsafe Thought 7.34% 78.11%
@@
Safe Thought 12.68% 2.74%
Unsafe Thought 25.83% 58.75%
@@
Safe Thought 15.89% 3.06%
Unsafe Thought 11.85% 69.23%
@@
Safe Thought 24.52% 6.91%
Unsafe Thought 6.19% 62.38%
@@
Safe Thought 19.88% 3.93%
Unsafe Thought 7.44% 68.75%
@@
Safe Thought 12.65% 1.90%
Unsafe Thought 7.34% 78.11%
R1-Onevision+Ours Safe Answer Unsafe Answer
Safe Thought 74.58% 8.87%
Unsafe Thought 16.55% 0.00%
@@
Safe Thought 15.89% 3.06%
Unsafe Thought 11.85% 69.23%
LLaVA-CoT+Ours Safe Answer Unsafe Answer
Safe Thought 71.73% 21.73%
Unsafe Thought 6.43% 0.06%
@@
Figure 7: Prompt used in data construction.
@@ D.3 Case Study of Ablation Experiments
Examples of responses generated by fine-tuned on TiS and TiS no thought dataset are illustrated in Figure 8 and Figure 9. The TiS dataset without thought is predominantly composed of unsafe examples. When fine-tuning MLRMs based on this data, the model tends to generate responses that consist solely of brief refusals. In particular, for safety-awareness examples, the model tends to either produce overly brief refusals or fail to detect unsafe risks. In contrary, the complete TiS data enables the detection of more unsafe intents during the thought process, while also providing more helpful information.
@@ D.4 Case Study
Examples of responses generated by fine-tuned on VLGuard, MIS, SPA-VL and TiS dataset are illustrated in Figure 10. Our approach demonstrates the ability to retain the thought process of the models while decisively rejecting unsafe inputs and explicitly articulating the potential serious consequences associated with such queries.

### v4

@@ 5 Related Work
With the rapid development and widespread deployment of LRMs, many works have paid attention to safety of LRMs. Several studies  Jiang et al. (2025); Parmar and Govindarajulu (2025) have conducted comprehensive safety evaluations on LRMs, e.g., DeepSeek-R1 Guo et al. (2025), revealing existing vulnerabilities of these models. In addition, Zhou et al. (2025) introduce a CoT jailbreak attack method, specifically targeting reasoning models. On the other hand, to mitigate the issues of safety, recent studies Jiang et al. (2025); Zhang et al. (2025b) have constructed the safety data with the thought process for supervised fine-tuning of LRMs. However, the above-mentioned works are all confined to the text-only models. Although SafeMLRM Fang et al. (2025) explores the safety of MLRMs and reveals three critical findings such as Reasoning Tax, the limitation of this work is that it only focuses on one safety scenario, i.e., jailbreak robustness. Our study is the first to investigate the safety behaviors of multimodal reasoning models in both jailbreak and awareness of safety scenarios and reveals more findings beyond Reasoning Tax.
@@
Current studies enhancing the safety capabilities of MLLMs can be categorized into two types: the training-based method and the training-free method. The training-based method typically includes Supervised Fine-Tuning (SFT) Zong et al. (2024); Ding et al. (2025b); Li et al. (2024) and Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2024). In addition, Lee et al. (2024) introduce the weight merging approach to mitigate safety degradation. Another branch of studies incorporates additional safety components in a training-free manner. For instance, MLLM-Protector Pi et al. (2024) offers a plug-and-play detector for harmful responses, ECSO Gou et al. (2024) converts the image inputs into text to exploit the safety of pre-aligned LLMs and ETA Ding et al. (2025a) proposes an inference-time alignment framework to ensure safety compliance. In contrast to these methods, our study is the first to propose potential solutions to address the issue of safety degradation of MLRMs.
@@ 6 Conclusion
Ethical Considerations
In this paper, we primarily focus on investigation on the safety evaluation of MLRMs. All experiments are conducted using publicly released datasets in a controlled setting, thereby avoiding the creation or propagation of new harmful content. We highlight that the goal of our work is to reveal severe safety essue exhibited by MLRMs. Moreover, we designed TiS dataset to support the development of safer MLRMs without raising ethical concerns.
Acknowledgments
The research work descried in this paper has been supported by the National Nature Science Foundation of China (No. 62376019, 62476023, 61976015, 61976016, 61876198 and 61370130), and the National Key R&D Program of China (2020AAA0108001). The authors would like to thank the anonymous reviewers for their valuable comments and suggestions to improve this paper.
@@ A.1 Benchmark and Evaluation Details
For the MM-SafetyBench, FigStep and SIUO, we employ GPT-4o-mini Hurst et al. (2024) and adopt identical configurations in their work to ensure consistency across assessments. For the JailBreaKV, we follow the same evaluation settings of their work and utilize the Llama Guard 3 Vision Chi et al. (2024) as the evaluator. However, similar to the issues observed in MIS Ding et al. (2025b), we identify shortcomings in the safe category evaluation setup of MSSBench. As a result, we employed human evaluation as an alternative. Additional details are provided in Appendix A.2. The benchmarks are listed as follows:
@@
JailBreaKV Luo et al. (2024) is designed to evaluate the transferability of jailbreak techniques from LLMs to MLLMs, encompassing both text-based and image-based jailbreak attacks. Following the setting of VLSBench Hu et al. (2024), we conduct our experiments on the miniset provided in their paper, which contains 280 samples.
@@ A.2 MSSBench Evaluation Details
In the evaluation setup of MSSBench, LLMs are required to classify responses into four safety levels. For certain examples in the safe category, responses that provide helpful advice but also emphasizing potential risks are evaluated as incorrect. As shown in Figure 7, the model provides the suggestion of using a heavy-duty drill while emphasizes safety as a priority. To address this limitation, we employed human evaluation as an alternative.
@@ A.3 Model Details
We evaluate a total of 13 MLRMs, Table 6 summarizes the Multimodal Large Reasoning Models evaluated for safety and their corresponding base models.
Model Chat Embodied AVG.
Safe Unsafe AVG. Safe Unsafe AVG.
[…]
@@ B.1 Detailed Pipeline
To accommodate the input requirements of DeepSeek-R1 Guo et al. (2025), we convert images into caption form as input in advance.In particular, we employ Qwen2.5-VL-72B Bai et al. (2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of safety issues, with the prompt “Please provide a detailed description of this image.”.
@@
The primary objective in this step is to filter and transform the original thought process. In particular, the caption field is processed to prevent potential misinterpretation by the MLRMs. The original thought process from the previous step also contains redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted. Specific prompts used for this step are provided in Figure 8.
@@ B.3 Prompt for Data Construction
This section presents the prompt and safety regulation used for safety thought process collection. We referenced the definition of User Requests Categorization from Wang et al. (2024) and listed all prompts used in the data construction in Figure 8. During the collection of safety thought process data, we supplied the model with category-specific safety regulations derived from the raw data classification. This strategy not only reduces the length of prompt but also enables tiered defensive measures for distinct safety-related issues.
@@ C.2 Traning Details
For the training procedure, we utilize the LoRA Hu et al. (2022) to fine-tune the R1-Onevision and LLaVA-CoT based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework. For the SPA-VL dataset, to ensure fairness, we randomly selected 4,000 data samples with balanced categories and used the DPO script for training. All training experiments are conducted using one NVIDIA A100-80G GPU. The hyper-parameters for supervised fine-tuning are listed in Table 9.
@@ D.2 Performance of Thought Process on MM-SafetyBench
Safe Thought 13.71% 41.66%
Unsafe Thought 6.72% 37.91%
@@
Safe Thought 13.27% 40.65%
Unsafe Thought 25.25% 20.83%
@@
Safe Thought 18.69% 40.77%
Unsafe Thought 9.05% 31.49%
@@
Safe Thought 23.15% 44.76%
Unsafe Thought 7.50% 24.52%
@@
Safe Thought 21.07% 42.26%
Unsafe Thought 6.25% 30.42%
@@
Safe Thought 13.71% 41.66%
Unsafe Thought 6.72% 37.91%
R1-Onevision+TiS Safe Answer Unsafe Answer
Safe Thought 77.32% 21.49%
Unsafe Thought 0.95% 0.24%
@@
Safe Thought 18.69% 40.77%
Unsafe Thought 9.05% 31.49%
LLaVA-CoT+TiS Safe Answer Unsafe Answer
Safe Thought 84.82% 8.75%
Unsafe Thought 5.89% 0.54%
@@
Figure 8: Prompts used in data construction.
@@ D.3 Case Study of Ablation Experiments
Examples of responses generated by models fine-tuned on TiS and TiS no thought dataset are illustrated in Figure 9 and Figure 10. The TiS dataset without thought is predominantly composed of unsafe examples. When fine-tuning MLRMs based on this data, the model tends to generate responses that consist solely of brief refusals. In particular, for safety-awareness examples, the model tends to either produce overly brief refusals or fail to detect unsafe risks. In contrast, the complete TiS data enables the detection of more unsafe intents during the thought process, while also providing more helpful information.
@@ D.4 Case Study
Examples of responses generated models fine-tuned on VLGuard, MIS, SPA-VL and TiS dataset are illustrated in Figure 11 and Figure 12. Our approach demonstrates the ability to retain the thought process of the models while decisively rejecting unsafe inputs and explicitly articulating the potential serious consequences associated with such queries.